# spacing=nonfrench

### v1

@@ 8.1. Methodology
Scaling servers. Although our testbed contains only two physical servers, our Tofino switch has significantly higher forwarding throughput (3.2 Tbps (swi, [n. d.])) than server throughput (tens of KOPS) and is not our evaluation bottleneck. To simulate larger-scale deployment with limited hardware while maintaining realism in evaluation, we adopt the well-known server rotation approach as in prior in-switch caching studies (Jin et al., 2017; Sheng et al., 2023). Let $N$ be the number of simulated servers. We assign files to $N$ simulated servers using the HDFS’s RBF policy (§7). We calculate each server’s load from file access frequencies, and identify the bottlenecked server with the highest load. We then perform our evaluation in $N$ iterations. We first deploy the bottlenecked server in one physical machine and fully saturate it to measure its performance. In the following N1N-1 iterations, we reset the physical machines to the initial state, pair the bottlenecked server with one of the N1N-1 non-bottlenecked servers, deploy them on the two physical machines, and measure the performance of the non-bottlenecked server. We aggregate the performance of all servers as the overall performance. By default, we simulate 16 servers and increase the scale to 128 servers (Exp#1).
@@
Baselines. We consider client-side caching (CCache) by faithfully following the state-of-the-art client-side caching implementations in IndexFS (Ren et al., 2014) and InfiniFS (Lv et al., 2022)9). CCache uses RocksDB (v6.22.1) as a key-value store to keep all metadata instead of in an HDFS namenode in each simulated server to eliminate HDFS’s path resolution overhead. Each CCache client locally caches directories’ permission metadata, and forwards all read requests to servers for attribute retrieval. We implement lazy invalidation (Lv et al., 2022) for cache consistency, which outperforms lease-based cache management (Ren et al., 2014). We do not compare FMCache with in-switch key-value caches (e.g., NetCache (Jin et al., 2017) and FarReach (Sheng et al., 2023)) since they are designed for key-value stores and do not support file-system semantics (i.e., they cannot ensure correctness and consistency for file-system operations).
@@
We evaluate four schemes: (i) NoCache, which does not employ any caching and performs metadata operations directly with HDFS namenodes in all simulated servers; (ii) CCache, our client-side caching implementation; (iii) FMCache, which extends NoCache with in-switch caching; and (iv) FMCache+, which extends CCache with in-switch caching.
@@
Before each experiment, we pre-load 32 million files into HDFS namenodes for NoCache and FMCache, and the corresponding paths and metadata into RocksDB for CCache and FMCache+. We pre-load the 5,000 hottest files and their ancestors into the in-switch cache for FMCache and FMCache+. We set the pre-defined CMS threshold of FMCache and FMCache+ to 10 and 20, respectively (a larger threshold is used for FMCache+ due to its higher throughput with client-side caching), and reset the CMS and the frequency counter array every two seconds. We simulate 128 client threads, which sufficiently saturate back-end servers. The client-side cache of each simulated client in CCache and FMCache+ is allocated 4 MiB (Lv et al., 2022). We plot the average results over five runs, with error bars as 95% confidence intervals under the Student’s t-distribution.
8.2. Performance Analysis
@@
Figure 7. (Exp#1) Performance under real-world workloads.
@@
(Exp#1) Throughput under real-world workloads. Figure 7 shows the throughput results on 16 and 128 simulated servers via server rotation under four real-world workloads. Under 16 servers, FMCache increases NoCache’s throughput by 22.9%, 51.7%, 53.7%, and 47.6% in Alibaba, Training, Thumb, and LinkedIn, respectively, due to load balancing via in-switch caching, while NoCache suffers from load imbalance.
@@
FMCache has lower throughput than CCache by around 38.1% for all workloads. The reason is that CCache reduces the path resolution overhead of HDFS by caching directories’ permission metadata in clients and bypasses the overhead of lease management and distributed transactions in HDFS. However, by integrating FMCache with CCache, FMCache+ increases CCache’s throughput by 8.0%, 33.5%, 36.3%, and 31.5% in Alibaba, Training, Thumb, and Linkedin, respectively, as CCache forwards all read requests to servers for attribute retrieval (§8.1) while FMCache+ improves load balancing via in-switch caching. Our results show that FMCache complements client-side caching to improve metadata performance.
@@
Under 128 servers, FMCache increases NoCache’s throughput by 11.0%, 134.6%, 181.6%, and 71.2% in Alibaba, Training, Thumb, and LinkedIn, respectively. FMCache+ increases CCache’s throughput by 14.7%, 103.6%, 139.6%, and 57.3% for the same workloads, respectively. FMCache and FMCache+ achieve significantly higher throughput gains than that in 16 servers (except in Alibaba) as they improve server scalability via load balancing. The throughput gains from 16 to 128 servers are marginal in Alibaba, which has the largest write ratio among the four workloads. Since FMCache and FMCache+ adopt write-through caching, maintaining cache consistency for extensive writes incurs substantial overhead.
@@
(Exp#2) Single-operation performance. Figure 8 shows the throughput of individual metadata operations: open, stat, create, mkdir, rename, chmod, delete, and rmdir. For read operations (open and stat), FMCache increases NoCache’s throughput by 80.1% and 80.5%, while FMCache+ increases CCache’s throughput by 74.9% and 77.5%, respectively. For write operations (create, mkdir, rename, chmod, delete, and rmdir), FMCache has lower throughput than NoCache by 2.7%, 0.2%, 13.2%, 36.5%, 12.7%, and 14.6%, respectively, while FMCache+ has lower throughput than CCache, by 4.9%, 3.5%, 7.2%, 12.3%, 8.1%, and 6.4%, respectively. The performance drops stem from the switch’s cache maintenance overhead. Among write operations, chmod incurs the highest overhead as chmod on cached paths requires fetching metadata from HDFS namenodes and updating the in-switch cache with the latest metadata. In contrast, rename, delete, and rmdir on cached paths only mark the paths as deleted in the in-switch cache, while create and mkdir only involve creations on new files and directories on HDFS namenodes, respectively, and will not trigger cache updates.
Figure 8. (Exp#2) Single-operation performance.
@@
(Exp#3) Impact of chmod ratio. We further analyze chmod, which triggers frequent cache updates and shows the most performance drops in FMCache and FMCache+ (Exp#2). We generate mixed read-write workloads composed of open (i.e., reads) and chmod (i.e., writes) with different ratios. Each of open and chmod follows a power-law distribution with an exponent of 0.9. Figure 9 shows that at 0% chmod ratio, FMCache and FMCache+ achieve higher throughput than NoCache and CCache, respectively, but as the chmod ratio increases, their throughput decreases, while the throughput of NoCache and CCache remains stable. FMCache and FMCache+ begin to show throughput degradations when the chmod ratio exceeds 50%. At 100% chmod ratio, FMCache and FMCache+ reach the lowest throughput, 36.5% and 12.3% lower than NoCache and CCache, respectively. Nevertheless, real-world workloads have low chmod ratios (Table 1), so FMCache and FMCache+ still maintain performance gains in practice (Exp#1).
@@
FMCache’s multi-level read-write locking design (§5) significantly reduces the overhead in writes. We compare multi-level locking (MultiLock) with single-level locking (SingleLock), which always maps a full path to the first lock counter array. We measure the recirculation count (i.e., number of recirculations per request) by the monitoring tool Barefoot Shell (bfs, [n. d.]). We do not compare MultiLock and locking-disabled in-switch caches, which cannot maintain cache consistency. Table 2 shows that for read-only and write-only workloads, the recirculation counts for both SingleLock and MultiLock are 7.63 and 1, respectively, as FMCache recirculates a read request multiple times for path resolution and a write request once for lock access. MultiLock significantly reduces the recirculation count of SingleLock (e.g., by 93.1% for 75% chmod) due to mitigated lock contention.
Figure 9. (Exp#3) Throughput versus chmod ratio.
@@
Table 2. (Exp#3) Recirculation count versus chmod ratio.
(Exp#4) Latency analysis. We analyze request latencies by adjusting the request sending rate to a target throughput, as in prior studies (Cheng et al., 2015; Jin et al., 2017; Sheng et al., 2023; Didona and Zwaenepoel, 2019), so as to analyze the trade-off between latencies and throughput. We focus on (i) a read-only workload that issues 32 million open requests and (ii) the Alibaba workload, which has the largest write ratio (Table 1). The two workloads show FMCache’s best- and worst-case performance, respectively, under write-through caching (§3.1). We follow a power-law distribution with an exponent of 0.9 and consider 16 simulated servers.
@@
Figure 10. (Exp#4) Latency analysis.
@@
Figure 10 shows the average, p95, and p99 latency results. For the read-only workload, all schemes show low latencies at low target throughput (less than 0.1 MOPS) as the servers are not saturated and do not have queueing delays. When the target throughput increases to 0.2 MOPS, NoCache starts to show increasing latencies due to load imbalance under higher loads. FMCache maintains low latencies via in-switch caching. For example, at target throughout 0.26 MOPS, FMCache reduces NoCache’s average, p95, and p99 latencies by 64.6%, 89.8%, and 87.7%, respectively. The same trend is also observed for FMCache+ and CCache. For example, at target throughput 0.59 MOPS, FMCache+ reduces CCache’s average, p95, and p99 latencies by 26.9%, 14.7%, and 77.0%, respectively.
@@
For the Alibaba workload, FMCache also outperforms NoCache. For example, at target throughput 0.15 MOPS, FMCache reduces NoCache’s average, p95 and p99 latencies by 66.1%, 65.4%, and 37.2%, respectively. For FMCache+ and CCache, at low throughput, they have comparable average and p95 latencies, while FMCache has slightly higher p99 latencies due to the switch’s cache maintenance overhead. At high throughput (e.g., 0.35 MOPS), FMCache+ reduces CCache’s average, p95, and p99 latencies by 25.5%, 63.7%, and 15.5%, respectively.
8.3. Impact of Workload Settings
We examine FMCache’s robustness across different workload settings. In the interest of space, we focus on Alibaba (with the largest write ratio) and Thumb (with the largest read ratio), while Training and LinkedIn show consistent behaviors with Thumb due to their similar read ratios.
@@ 8.3. Impact of Workload Settings
(Exp#5) Impact of file access frequency assignment. We generate a sequence of access frequencies in descending order based on a power-law distribution with an exponent of 0.9, and another sequence of files based on different sorting orders. We assign the $i$-th access frequency to the $i$-th file. We consider three sorted file sequence: (i) high-level-first (HLF), which sorts files in descending order of their levels (i.e., files at higher levels have higher access frequencies), (ii) low-level-first (LLF), which sorts files in ascending order of their levels (i.e., files at lower levels have higher access frequencies), and (iii) random (our default), which generates a random sequence of files across all levels. Figure 11 shows that FMCache and FMCache+ still outperform NoCache and CCache, respectively, under different file access frequency assignments. For example, in Thumb (the most read-intensive), FMCache and FMCache+ increase the throughput of NoCache and CCache by 53.7-69.0% and 30.2-44.7%, respectively.
@@
(Exp#6) Impact of access skewness. We vary the skewness level for access frequencies under the uniform distribution and varying power-law distributions with an exponent of 0.8, 0.9 (our default), and 1.0. A larger exponent implies a more skewed access pattern. Figure 12 shows that under the uniform workload, FMCache and FMCache+ have slightly less throughput than NoCache and CCache, respectively, by up to 5.0%, since most requests come from uncached paths and are served by servers, while FMCache and FMCache+ incur extra cache maintenance overhead.
@@
(b) Thumb
Figure 11. (Exp#5) Impact of file access frequency assignment.
@@
(b) Thumb
Figure 12. (Exp#6) Impact of access skewness.
@@
(b) Thumb
Figure 13. (Exp#7) Impact of maximum path depth.
[…]
For Alibaba, as the exponent increases from 0.9 to 1.0, the throughput of FMCache and FMCache+ decreases. The reason is that Alibaba has the largest write ratio and incurs significant cache maintenance overhead. Nevertheless, at exponent 1.0, FMCache and FMCache+ still increase the throughput of NoCache and CCache by 37.6% and 25.5%, respectively.
(Exp#7) Impact of maximum path depth. We vary the maximum path depth as 3, 5, 7, and 9 (our default). Figure 13 shows that FMCache and FMCache+ always outperform NoCache and CCache, respectively. For example, in Thumb, FMCache and FMCache+ increase the throughput of NoCache and CCache by 53.7-77.9% and 34.1-36.3%, respectively.
(Exp#8) Impact of dynamic workloads. We consider dynamic workloads with varying access frequencies of files over time. We follow the prior studies (Li et al., 2016; Jin et al., 2017; Sheng et al., 2023) to generate the hot-in dynamic pattern, which periodically selects the 100 least-frequently accessed files, re-assigns them with the highest access frequencies, and adjusts the access frequencies of other files accordingly to maintain a power-law distribution. We set the change period as 20 seconds and run each scheme for 200 seconds to measure per-second throughput. We disable server rotation as the system states change under dynamic workloads; instead, we issue workloads to the two physical servers and measure performance directly.
Figure 14 shows that for Thumb, FMCache and FMCache+ show performance dips due to periodic changes of file access frequencies. Before new hot records are admitted, performance dips occur, but FMCache and FMCache+ quickly admit new hot records and return to high performance with path-aware cache management (§4). Also, local hash collision resolution (§6) incurs minimal overhead to cache admission and eviction. For Alibaba, which has the largest write ratio, FMCache and FMCache+ have marginal performance gains over NoCache and CCache, respectively. Nevertheless, FMCache and FMCache+ still effectively respond to dynamic workloads.
@@
(b) Thumb
Figure 14. (Exp#8) Impact of dynamic workloads.
8.4. Switch Deployment
(Exp#9) Switch resource usage. We measure the resource usage on the Tofino switch for (i) SRAM (15 MiB), (ii) number of stages (12), (iii) number of ALUs (48), and (iv) PHV size (768 bytes). We also quote the resource usage of two in-switch key-value caches, NetCache (Jin et al., 2017) and FarReach (Sheng et al., 2023). Table 3 shows that NoCache and CCache use the least resources for L2/L3 forwarding, while FMCache and FMCache+ also support L2/L3 forwarding to process packets unrelated to metadata operations at line rate (3.2 Tbps for the Tofino switch (swi, [n. d.])). Additionally, FMCache and FMCache+ consume additional resources for caching (e.g., using SRAM and ALUs to cache metadata, track access frequencies, and maintain lock counters, and using PHVs to parse metadata requests), yet their resource usage is comparable to NetCache and FarReach (state-of-the-art in-switch key-value caches).

### v2

@@ VIII-A Methodology
Scaling servers. Although our testbed contains only two physical servers, our Tofino switch has significantly higher forwarding throughput (3.2 Tbps [switchthpt]) than server throughput (tens of KOPS) and is not our evaluation bottleneck. To simulate larger-scale deployment with limited hardware while maintaining realism in evaluation, we adopt the well-known server rotation approach as in prior in-switch caching studies [jin17, sheng25]. Let $N$ be the number of simulated servers. We assign files to $N$ simulated servers using the HDFS’s RBF policy (§VII). We calculate each server’s load from file access frequencies, and identify the bottlenecked server with the highest load. We then perform our evaluation in $N$ iterations. We first deploy the bottlenecked server in one physical machine and fully saturate it to measure its performance. In the following N1N-1 iterations, we reset the physical machines to the initial state, pair the bottlenecked server with one of the N1N-1 non-bottlenecked servers, deploy them on the two physical machines, and measure the performance of the non-bottlenecked server. We aggregate the performance of all servers as the overall performance. By default, we simulate 16 servers and increase the scale to 128 servers (Exp#1).
@@
Baselines. We consider client-side caching (CCache) by faithfully following the state-of-the-art client-side caching implementations in IndexFS [ren14] and InfiniFS [lv22]IX). CCache uses RocksDB (v6.22.1) as a key-value store to keep all metadata instead of in an HDFS namenode in each simulated server to eliminate HDFS’s path resolution overhead. Each CCache client locally caches directories’ permission metadata, and forwards all read requests to servers for attribute retrieval. We implement lazy invalidation [lv22] for cache consistency, which outperforms lease-based cache management [ren14]. We do not compare Fletch with in-switch key-value cache systems (e.g., NetCache [jin17] and FarReach [sheng25]) since they are designed for key-value stores and do not support file-system semantics (i.e., they cannot ensure correctness and consistency for file-system operations).
@@
We evaluate four schemes: (i) NoCache, which does not employ any caching and performs metadata operations directly with HDFS namenodes in all simulated servers; (ii) CCache, our client-side caching implementation; (iii) Fletch, which extends NoCache with in-switch caching; and (iv) Fletch+, which extends CCache with in-switch caching.
@@
Before each experiment, we pre-load 32 million files into HDFS namenodes for NoCache and Fletch, and the corresponding paths and metadata into RocksDB for CCache and Fletch+. We pre-load the 5,000 hottest files and their ancestors into the in-switch cache for Fletch and Fletch+. We set the pre-defined CMS threshold of Fletch and Fletch+ to 10 and 20, respectively (a larger threshold is used for Fletch+ due to its higher throughput with client-side caching), and reset the CMS and the frequency counter array every two seconds. We simulate 128 client threads, which sufficiently saturate back-end servers. The client-side cache of each simulated client in CCache and Fletch+ is allocated 4 MiB [lv22]. We plot the average results over five runs, with error bars as 95% confidence intervals under the Student’s t-distribution.
VIII-B Performance Analysis
@@
Figure 7: (Exp#1) Performance under real-world workloads.
@@
(Exp#1) Throughput under real-world workloads. Figure 7 shows the throughput results on 16 and 128 simulated servers via server rotation under four real-world workloads. Under 16 servers, Fletch increases NoCache’s throughput by 22.9%, 51.7%, 53.7%, and 47.6% in Alibaba, Training, Thumb, and LinkedIn, respectively, due to load balancing via in-switch caching, while NoCache suffers from load imbalance.
@@
Fletch has lower throughput than CCache by around 38.1% for all workloads. The reason is that CCache reduces the path resolution overhead of HDFS by caching directories’ permission metadata in clients and bypasses the overhead of lease management and distributed transactions in HDFS. However, by integrating Fletch with CCache, Fletch+ increases CCache’s throughput by 8.0%, 33.5%, 36.3%, and 31.5% in Alibaba, Training, Thumb, and Linkedin, respectively, as CCache forwards all read requests to servers for attribute retrieval (§VIII-A) while Fletch+ improves load balancing via in-switch caching. Our results show that Fletch complements client-side caching to improve metadata performance.
@@
Under 128 servers, Fletch increases NoCache’s throughput by 11.0%, 134.6%, 181.6%, and 71.2% in Alibaba, Training, Thumb, and LinkedIn, respectively. Fletch+ increases CCache’s throughput by 14.7%, 103.6%, 139.6%, and 57.3% for the same workloads, respectively. Fletch and Fletch+ achieve significantly higher throughput gains than that in 16 servers (except in Alibaba) as they improve server scalability via load balancing. The throughput gains from 16 to 128 servers are marginal in Alibaba, which has the largest write ratio among the four workloads. Since Fletch and Fletch+ adopt write-through caching, maintaining cache consistency for extensive writes incurs substantial overhead.
@@
(Exp#2) Single-operation performance. Figure 8 shows the throughput of individual metadata operations: open, stat, create, mkdir, rename, chmod, delete, and rmdir. For read operations (open and stat), Fletch increases NoCache’s throughput by 80.1% and 80.5%, while Fletch+ increases CCache’s throughput by 74.9% and 77.5%, respectively. For write operations (create, mkdir, rename, chmod, delete, and rmdir), Fletch has lower throughput than NoCache by 2.7%, 0.2%, 13.2%, 36.5%, 12.7%, and 14.6%, respectively, while Fletch+ has lower throughput than CCache, by 4.9%, 3.5%, 7.2%, 12.3%, 8.1%, and 6.4%, respectively. The performance drops stem from the switch’s cache maintenance overhead. Among write operations, chmod incurs the highest overhead as chmod on cached paths requires fetching metadata from HDFS namenodes and updating the in-switch cache with the latest metadata. In contrast, rename, delete, and rmdir on cached paths only mark the paths as deleted in the in-switch cache, while create and mkdir only involve creations on new files and directories on HDFS namenodes, respectively, and will not trigger cache updates.
Figure 8: (Exp#2) Single-operation performance.
@@
(Exp#3) Impact of chmod ratio. We further analyze chmod, which triggers frequent cache updates and shows the most performance drops in Fletch and Fletch+ (Exp#2). We generate mixed read-write workloads composed of open (i.e., reads) and chmod (i.e., writes) with different ratios. Each of open and chmod follows a power-law distribution with an exponent of 0.9. Figure 9 shows that at 0% chmod ratio, Fletch and Fletch+ achieve higher throughput than NoCache and CCache, respectively, but as the chmod ratio increases, their throughput decreases, while the throughput of NoCache and CCache remains stable. Fletch and Fletch+ begin to show throughput degradations when the chmod ratio exceeds 50%. At 100% chmod ratio, Fletch and Fletch+ reach the lowest throughput, 36.5% and 12.3% lower than NoCache and CCache, respectively. Nevertheless, real-world workloads have low chmod ratios (Table I), so Fletch and Fletch+ still maintain performance gains in practice (Exp#1).
@@
Fletch’s multi-level read-write locking design (§V) significantly reduces the overhead in writes. We compare multi-level locking (MultiLock) with single-level locking (SingleLock), which always maps a full path to the first lock counter array. We measure the recirculation count (i.e., number of recirculations per request) by the monitoring tool Barefoot Shell [bfshell]. We do not compare MultiLock and locking-disabled in-switch caches, which cannot maintain cache consistency. Table II shows that for read-only and write-only workloads, the recirculation counts for both SingleLock and MultiLock are 7.63 and 1, respectively, as Fletch recirculates a read request multiple times for path resolution and a write request once for lock access. MultiLock significantly reduces the recirculation count of SingleLock (e.g., by 93.1% for 75% chmod) due to mitigated lock contention.
Figure 9: (Exp#3) Throughput versus chmod ratio.
Table II: (Exp#3) Recirculation count versus chmod ratio.
@@
(Exp#4) Latency analysis. We analyze request latencies by adjusting the request sending rate to a target throughput, as in prior studies [cheng15, jin17, sheng25, didona19], so as to analyze the trade-off between latencies and throughput. We focus on (i) a read-only workload that issues 32 million open requests and (ii) the Alibaba workload, which has the largest write ratio (Table I). The two workloads show Fletch’s best- and worst-case performance, respectively, under write-through caching (§III-A). We follow a power-law distribution with an exponent of 0.9 and consider 16 simulated servers.
@@
Figure 10: (Exp#4) Latency analysis.
@@
Figure 10 shows the average, p95, and p99 latency results. For the read-only workload, all schemes show low latencies at low target throughput (less than 0.1 MOPS) as the servers are not saturated and do not have queueing delays. When the target throughput increases to 0.2 MOPS, NoCache starts to show increasing latencies due to load imbalance under higher loads. Fletch maintains low latencies via in-switch caching. For example, at target throughout 0.26 MOPS, Fletch reduces NoCache’s average, p95, and p99 latencies by 64.6%, 89.8%, and 87.7%, respectively. The same trend is also observed for Fletch+ and CCache. For example, at target throughput 0.59 MOPS, Fletch+ reduces CCache’s average, p95, and p99 latencies by 26.9%, 14.7%, and 77.0%, respectively.
@@
For the Alibaba workload, Fletch also outperforms NoCache. For example, at target throughput 0.15 MOPS, Fletch reduces NoCache’s average, p95 and p99 latencies by 66.1%, 65.4%, and 37.2%, respectively. For Fletch+ and CCache, at low throughput, they have comparable average and p95 latencies, while Fletch has slightly higher p99 latencies due to the switch’s cache maintenance overhead. At high throughput (e.g., 0.35 MOPS), Fletch+ reduces CCache’s average, p95, and p99 latencies by 25.5%, 63.7%, and 15.5%, respectively.
VIII-C Impact of Workload Settings
@@
(Exp#5) Impact of file access frequency assignment. We generate a sequence of access frequencies in descending order based on a power-law distribution with an exponent of 0.9, and another sequence of files based on different sorting orders. We assign the $i$-th access frequency to the $i$-th file. We consider three sorted file sequence: (i) high-level-first (HLF), which sorts files in descending order of their levels (i.e., files at higher levels have higher access frequencies), (ii) low-level-first (LLF), which sorts files in ascending order of their levels (i.e., files at lower levels have higher access frequencies), and (iii) random (our default), which generates a random sequence of files across all levels. Figure 11 shows that Fletch and Fletch+ still outperform NoCache and CCache, respectively, under different file access frequency assignments. For example, in Thumb (the most read-intensive), Fletch and Fletch+ increase the throughput of NoCache and CCache by 53.7-69.0% and 30.2-44.7%, respectively.
@@
(Exp#6) Impact of access skewness. We vary the skewness level for access frequencies under the uniform distribution and varying power-law distributions with an exponent of 0.8, 0.9 (our default), and 1.0. A larger exponent implies a more skewed access pattern. Figure 12 shows that under the uniform workload, Fletch and Fletch+ have slightly less throughput than NoCache and CCache, respectively, by up to 5.0%, since most requests come from uncached paths and are served by servers, while Fletch and Fletch+ incur extra overhead for maintaining cache consistency.
@@
(b) Training
Refer to caption Refer to caption
(c) Thumb
(d) LinkedIn
Figure 11: (Exp#5) Impact of file access frequency assignment.
@@
(b) Training
Refer to caption Refer to caption
(c) Thumb
(d) LinkedIn
Figure 12: (Exp#6) Impact of access skewness.
[…]
For Alibaba and LinkedIn, as the exponent increases from 0.9 to 1.0, the throughput of Fletch and Fletch+ decreases. Alibaba has the largest write ratio, while LinkedIn has the highest chmod ratio. They incur significant maintenance overhead for cache consistency. Nevertheless, at exponent 1.0 for LinkedIn, Fletch and Fletch+ still increase the throughput of NoCache and CCache by 84.3% and 45.1%, respectively.
(Exp#7) Impact of maximum path depth. We vary the maximum path depth as 3, 5, 7, and 9 (our default). Figure 13 shows that Fletch and Fletch+ always outperform NoCache and CCache, respectively. For example, in Thumb, Fletch and Fletch+ increase the throughput of NoCache and CCache by 53.7-77.9% and 34.1-36.3%, respectively.
@@
(b) Training
Refer to caption Refer to caption
(c) Thumb
(d) LinkedIn
Figure 13: (Exp#7) Impact of maximum path depth.
(Exp#8) Impact of dynamic workloads. We consider dynamic workloads with varying access frequencies of files over time. We follow the prior studies [li16, jin17, sheng25] to generate the hot-in dynamic pattern, which periodically selects the 100 least-frequently accessed files, re-assigns them with the highest access frequencies, and adjusts the access frequencies of other files accordingly to maintain a power-law distribution. We set the change period as 20 seconds and run each scheme for 200 seconds to measure per-second throughput. We disable server rotation as the system states change under dynamic workloads; instead, we issue workloads to the two physical servers and measure performance directly.
Figure 14 shows that for Training, Thumb, and LinkedIn, Fletch and Fletch+ show performance dips due to periodic changes of file access frequencies. Before new hot records are admitted, performance dips occur, but Fletch and Fletch+ quickly admit new hot records and return to high performance with path-aware cache management (§IV). Also, local hash collision resolution (§VI) incurs minimal overhead to cache admission and eviction. For Alibaba, which has the largest write ratio, Fletch and Fletch+ have marginal performance gains over NoCache and CCache, respectively, as they incur cache consistency overhead. Nevertheless, Fletch and Fletch+ still effectively respond to dynamic workloads.
@@
(b) Training
Refer to caption Refer to caption
(c) Thumb
(d) LinkedIn
Figure 14: (Exp#8) Impact of dynamic workloads.
VIII-D Switch Deployment
(Exp#9) Switch resource usage. We measure the resource usage on the Tofino switch for (i) SRAM (15 MiB), (ii) number of stages (12), (iii) number of ALUs (48), and (iv) PHV size (768 bytes). We also quote the resource usage of two in-switch key-value cache systems, NetCache [jin17] and FarReach [sheng25], based on the egress switch resource usage reported in the paper [sheng25]. Table III shows that NoCache and CCache use the least resources for L2/L3 forwarding, while Fletch and Fletch+ also support L2/L3 forwarding to process packets unrelated to metadata operations at line rate (3.2 Tbps for the Tofino switch [switchthpt]). Additionally, Fletch and Fletch+ consume additional resources for caching (e.g., using SRAM and ALUs to cache metadata, track access frequencies, and maintain lock counters, and using PHVs to parse metadata requests), yet their resource usage is comparable to NetCache and FarReach (state-of-the-art in-switch key-value cache approaches).
Table III: (Exp#9) Switch resource usage (numbers in brackets refer to the fractions of used resources over total available ones). Numbers of NetCache and FarReach are quoted from [sheng25].